OpenStudy (ledah):

Write the equation of the line given the slope and the y-intercept. m = 2, y-intercept (0, -7)

OpenStudy (jit4won):

m = y- y1/x-x1

OpenStudy (anonymous):

y-yo=m(x-x0) y-(-7)=2(x-0) y+7=2x

OpenStudy (jit4won):

x1= 0 y1 = -7